You invested $8000, part at 5% and part at 6%. The total interest earned in one year what $468. How much was invested at each rate?
Answer by checkley77(12569)

.06x+.05(8,000-x)=468
.06x+400-.05x=468
.01x=468-400
.01x=68
x=68/.01
x=6,800 amount invested @ 6%.
8,000-6,800=1,200 amount invested @ 5%.
Proof:
.06*6,800+.05*1,200=468
408+60=468
468=468